OpenStudy (anonymous):

solve for x? 6^-x-2=13^-8x

OpenStudy (experimentx):

take log on both sides,

OpenStudy (anonymous):

$6^{-x-2}=13^{-8x}$ ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[6^{-x-2}=13^{-8x}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[\log_{10}{6^{-x-2}}= \log_{10}{13^{-8x\ [\rightarrow x = (-2\log_{10}{6} )/(\log_{10}{6}-8\log_{10}{} )\]}}\]

OpenStudy (anonymous):

\[6^{-x-2}=13^{-8x}\implies (x+2)\ln(6)=8xln(13)\implies 8x \ln(13)-x \ln(6)=2 \ln(6)\]\[=x(8 \ln(13)- \ln(6))=2 \ln(6)\implies x=\frac{2 \ln(6)}{8 \ln(13)- \ln(6)}\]

OpenStudy (anonymous):

the same log 10 or log e

OpenStudy (anonymous):

Pretty cool, no?

OpenStudy (anonymous):

ok, cool

OpenStudy (anonymous):

Yours was perfectly correct, but I thought maybe somebody would benefit from a little more detail.

OpenStudy (anonymous):

yea, i expalin no detail